Preliminary Agenda for the VISTA IOT on Jan 27, 2011 ==================================================== 1. VISTA general status and planned activities (TSz) 2. Status of the surveys (MRe) 3. Status of various software: SADT/OT/P2PP/etc (MRe, MHi, JEm) 4. New/recent QC issues, data reduction/product issues (WHu, MIr) 5. Action Items review (enclosed bellow is the AI section from the minutes of the last IOT) 6. Other idssues 7. Date of the next meeting, proposed Mar 10-15 or Apr 1-10, 2011 - to avoid the proposal deadline rush. ************************************************************ ====================================================== ======= 6. Action Items review ============ ====================================================== OLD AIs: -------- AI 2009-11: TSz should find out if the constraints sets could be carried with the OBs, to the OS and the fits header of the VIRCAM files. Otherwise it is not easy to propagate these data by interfaces. 2009-07-29: TSz iterating with R. Schmutzer, but DFI (T. Bierwirth) will also need to be involved. For surveys, given the small range of programs it is expected that there are not too many different OBs, and this might need to be followed in a different/manual way in the beginning - this could imply a major change for the tools that cannot be done in the last moment. 2010-03-23: the keywords have been implemented except for the moon; there is a ticket with details; some testing done by VDI, no errors found; keep ongoing, try to complete during my next shift end of Mar - early Apr. 2010-04-23: VDI reports one formula is missing 2010-05-27: Corresponding PPRS is still pending. VDI asks to keep this AI open until the PPRS is closed. 2010-06-29: No changes with respect to last meeting. VDI asks to keep this AI kept open/pending until the PPRS is closed. 2010-08-24: VDI not present at the meeting. It is not known if there was any progress. Therefore decided to keep this pending while waiting for final status from VDI. 2010-10-05: VDI: still pending, will force the issue with the SW during my next shift from Oct 19 to 30, 2010 STATUS: PENDING ** AI 2010-06: TSz+VDI to improve the TIO training, to prepare an operational manual 2010-05-27: Operations wiki has been updated. More training needed. 2010-06-29: Not much was done due to problems with IQ + intervention + work on IP. Keep the AI open. 2010-08-24: SMI comments that the training is an ongoing task. Therefore it seems that this could be closed. It is not clear about the status of the operational manual mentioned in the AI. Thus for the moment kept as PENDING. Comment from VDI and TSZ on the status required. 2010-10-05: to be kept open for now as a reminder STATUS: PENDING ** AI 2010-09: photometric zeropoints as calculated by the pipeline. 2010-05-27: Ongoing, see corresponding discussion this meeting 2010-06-29: TSZ points out that this is not well defined AI. MIR adds that the currently existing plots already would have shown the problems with zero points from pawprints to tiles. It is not clear whether this is about the zero points from Garching pipeline? WHU states that Garching pipeline has de-blending on. MIR: There is no significant increase in the scatter with respect to what one would expect from the random noise. It is agreed in the end to send one example data set and then the comparison should be done directly between Garching and CASU. 2010-08-24: This is still pending. WHU should send some example to JLE/MIR in order to do the checks and clarify what the problem actually is about. 2010-10-05: the example data were sent to MIr and JLe; the Cambridge pipeline results for the individual detector were as expected from "rms" noise considerations, though offset by 0.1 mag from the Garching results. JLe and WHu are attempting to identify the cause of this difference which is most likely due to differences in calibration frames used. JLe sent a new linearity curve to use; WHu reprocessed the data with the new linearity correction and there were no changes; JLe asked if he could get the calibration frames used from WHu. CASU use the "average" of all detector zero-points to monitor extinction/throughput. STATUS: PENDING ** AI 2010-12: VDI put in the User manual persistence amplitude and slope. Add to the instrument page the persistence report. 2010-08-24: MRE reports that the persistence is mentioned in the latest version of the manual, but not the amplitude and slope. The persistence report still needs to be added to the instrument web page (link in the news?). 2010-10-05: still pending, to be done during my next shift which is for VISTA, Oct 19-30, 2010. STATUS: PENDING ** AI 2010-14: SMI to raise the issue of degradation of coating (based on degradation of photometric zero points). The next coating should be scheduled. 2010-10-05: Serge G. said that this is pending, some tests are still pending, M2 is much worse than M1; studying the experience of GEMINI; M2 hasn't been recoated for ~3 yrs; no way to measure the mirror reflectivity separately for M1 and M2 without a reflectivity measuring device; M2 seems to have a film of gray material, while M1 has only spots at the edges MIr: I had a quick look, Z is affected most, almost linear in terms of mag/month!! VDI: The AI will be kept open as a reminder and it is becoming urgent. STATUS: PENDING